# Curriculum Direct Preference Optimization for Diffusion and Consistency Models

### v2

@@ 4 Experiments
Table 1: Text alignment, aesthetic and human preference scores obtained by the baseline (pre-trained) LCM and SD models versus the three fine-tuning strategies: DDPO, DPO and Curriculum DPO. The results represent averages over three runs. The best scores are highlighted in bold.
LCM - 0.7243 $\pm$ 0.0048 6.0490 $\pm$ 0.0162 0.2912 $\pm$ 0.0021
DDPO [6] 0.7490 $\pm$ 0.0036 6.3730 $\pm$ 0.0130 0.2952 $\pm$ 0.0011
DPO [70] 0.7502 $\pm$ 0.0045 6.4741 $\pm$ 0.0095 0.2990 $\pm$ 0.0010
Curriculum DPO (ours) 0.7548 $\pm$ 0.0041 6.6417 $\pm$ 0.0083 0.3237 $\pm$ 0.0012
SD - 0.6804 $\pm$ 0.0052 5.5152 $\pm$ 0.0166 0.2784 $\pm$ 0.0015
DDPO [6] 0.7629 $\pm$ 0.0040 5.8183 $\pm$ 0.0129 0.2854 $\pm$ 0.0012
DPO [70] 0.7614 $\pm$ 0.0049 5.7146 $\pm$ 0.0162 0.2827 $\pm$ 0.0008
Curriculum DPO (ours) 0.7703 $\pm$ 0.0036 5.8232 $\pm$ 0.0197 0.2856 $\pm$ 0.0015
@@
Dataset. To create the training datasets for the text alignment, aesthetics and human preference tasks, we use the same strategy as Black et al. [6]. Thus, we generate 500 images for each text prompt. For the text alignment and human preference tasks, the text prompts are based on the subject-verb-object (SVO) pattern, e.g. “a dog riding a bike”. Following Black et al. [6], we use a list of 45 animals and 3 types of activities. For the aesthetics task, the prompts are based on the “a photo of <an object>” template, where the object represents one of the same 45 animal classes. For each of the three tasks, there are 67,500 (prompt, image) pairs for training, and 6,750 pairs for evaluation.
@@
Generative models. We conduct our experiments by using two pre-trained text-to-image generative models: Stable Diffusion (SD) v1.5 [49] and Latent Consistency Model (LCM) [34]. The LCM checkpoint is trained using consistency distillation from a SD v1.5 checkpoint. To generate images with SD, we use the DDIM sampler with 50 steps. For LCM, we generate data using 8 steps of the Multistep Latent Consistency Sampling [34] procedure. Finally, we consider two different resolutions, $256\times 256$ pixels for SD and $768\times 768$ pixels for LCM, respectively.
@@
Reward models. To assess the alignment between a text prompt and a generated image, we use the cosine similarity to compare the embeddings produced by Sentence-BERT [48] for the original prompt and the caption generated by LLaVA [31]. To evaluate visual appeal (aesthetics), we use the LAION Aesthetics Predictor [55], a linear model applied on CLIP [45] that is trained on human-rated images and predicts a score on a scale from $1$ to $10$. Finally, for human preference estimation, we employ HPSv2 [73], a CLIP model fine-tuned on a dataset of image pairs ranked by humans.
@@
Training setup. We conduct all experiments on a single A100 GPU, training each model for 10,000 training iterations, with a batch size of 16 and two steps of gradient accumulation. The training time was approximately 2 days for each LCM experiment, while using 64GB of VRAM. The SD experiments took approximately 1 day each, while using 36 GB of VRAM. We employed the AdamW optimizer with a constant learning rate of $3\cdot 10^{-4}$. LoRA is activated in all experiments. For LCM, we utilize trainable low-rank matrices with a dimension of 64 and an equivalent $\alpha$. In the case of SD, we employ a rank of 8 and an $\alpha$ of 32. Following Wallace et al. [70], we set $\beta=5000$ for Diffusion-DPO. For Consistency-DPO, we set $\beta=200$ via grid search over the set $\{50,100,200,300,500\}$. Similarly, we use grid search over the set $\{3,5,7\}$ to determine the number of curriculum batches for Curriculum DPO. The optimal value is $B=5$. The number of training iterations per batch is set to $K=400$ for the first four batches, i.e. $H_{i}=K,\forall i\in\{1,2,3,4\}$. For the last batch, the number of iterations is set to $H_{5}=10000-4\cdot K$. Hence, the total number of training iterations is the same as for the baseline, i.e. $\sum_{i=1}^{B}H_{i}=10000$. Since all experiments involve reward models, the image pairs are sampled based on their preference score difference rather than their ranking difference, thus sidestepping the management of skewed preference score distributions.
@@
Results. For the quantitative assessment, we measure the reward scores after fine-tuning the models on each of the three tasks, corresponding to the three reward models. The results are shown in Table 1. As anticipated, the original models yield the lowest scores, indicating that fine-tuning is generally useful. Both DPO and DDPO exhibit better performance than the baseline, but without a clear winner among the two approaches. In contrast, Curriculum DPO is the most effective method, achieving the best reward score on each task. Moreover, the gains of Curriculum DPO are significant in most cases.
Figure 2: Qualitative results before and after fine-tuning for the text alignment task. The fine-tuning alternatives are: DDPO, DPO and Curriculum DPO. Best viewed in color.
@@
We present qualitative results for the text alignment task in Figure 2. Considering the examples from the first column, we observe that Curriculum DPO is the only method which generates the beetle and the bike in a coherent picture. Moreover, in the third column, our method is the only one able to generate the butterfly without additional humans around. Regarding the examples generated with SD, Curriculum DPO exhibits clear improvements over the other methods when generating the goose (last column) and the squirrel riding a bike (sixth column). Qualitative results for the aesthetics and human preference tasks are presented in Appendix A.3.
Table 2: Average ratings of the baseline models and the three fine-tuning strategies (DDPO, DPO and Curriculum DPO) based on our human evaluation study. Each human annotated 1,280 randomly sampled images with a rating from 1 to 5. The study was completed by 4 human evaluators. The best average ratings are highlighted in bold.
@@
- 2.997 $\pm$ 0.274 2.438 $\pm$ 0.162 3.184 $\pm$ 0.438 3.136 $\pm$ 0.367
DDPO [6] 2.878 $\pm$ 0.285 3.199 $\pm$ 0.165 2.819 $\pm$ 0.829 3.097 $\pm$ 0.463
DPO [70] 3.041 $\pm$ 0.257 3.038 $\pm$ 0.268 2.794 $\pm$ 0.301 3.056 $\pm$ 0.412
Curriculum DPO (ours) 3.642 $\pm$ 0.373 3.394 $\pm$ 0.287 3.244 $\pm$ 0.587 3.384 $\pm$ 0.559
Subjective human evaluation study. We conducted a human evaluation study to evaluate the prompt alignment and aesthetics of the two pre-trained models, SD and LCM, along with their fine-tuned versions based on DDPO, DPO and Curriculum DPO, respectively. We asked the annotators to rate each image with a grade from $1$ to $5$, considering the task-specific evaluation criteria (prompt alignment / aesthetics). For each prompt, the generated images were shuffled and presented in a random order to prevent any form of cheating from the annotators. More details about the annotation process are provided in Appendix A.4. We present the average ratings in Table 2. We observe that Curriculum DPO outperforms the other methods by considerable margins, regardless of the task.
@@
Figure 3: Ablation results obtained by varying the hyperparameter $\beta$ for Consistency-DPO, the number of training iterations per batch $K$ for Curriculum DPO, and number of batches $B$ for Curriculum DPO. Fine-tuned LCM models are compared with the pre-trained LCM baseline on the human preference task (where scores are given by the HPSv2 reward model).
Ablation studies. We conduct several ablation studies to explore the impact of various hyperparameters associated with Consistency-DPO and Curriculum DPO. In Figure 3(a), we show various values for $\beta$ and their impact on DPO when applied on consistency models. For values of $\beta$ higher or equal to 200, Consistency-DPO surpasses the pre-trained LCM baseline. In Figure 3(b), we present an ablation on the impact of $K$ (the number of training iterations per curriculum batch) on the human preference scores returned by HPSv2. We obtain better results for $K=300$ and $K=400$, although Curriculum DPO surpasses the baseline LCM for all values of $K$. In Figure 3(c), we present the HPSv2 score for various choices of the number of batches, $B$. Once again, Curriculum DPO consistently surpasses the pre-trained LCM, regardless of the number of training batches.
In addition to our exploration of hyperparameters, we present a detailed study of the key components of Curriculum DPO in Table 3. This analysis showcases the impact of the LoRA fine-tuning technique, the DPO objective, and the curriculum strategy. Interestingly, applying LoRA to LCM degrades performance on all three tasks. However, combining LoRA with DPO or Curriculum DPO leads to significantly better results. Curriculum DPO consistently outperforms its ablated version, justifying the proposed design.
Table 3: Ablation study on the key components of Curriculum DPO. LoRA, DPO and curriculum learning are gradually added to produce Curriculum DPO. The results represent averages over three runs. The best scores are highlighted in bold.
@@
Model LoRA DPO Curriculum Text Aesthetics Human
DPO Alignment Preference
LCM 0.7243 $\pm$ 0.0048 6.0490 $\pm$ 0.0162 0.2912 $\pm$ 0.0021
0.7151 $\pm$ 0.0032 5.7158 $\pm$ 0.0154 0.2881 $\pm$ 0.0017
0.7502 $\pm$ 0.0045 6.4741 $\pm$ 0.0095 0.2990 $\pm$ 0.0010
0.7548 $\pm$ 0.0041 6.6417 $\pm$ 0.0083 0.3237 $\pm$ 0.0012
SD 0.6804 $\pm$ 0.0052 5.5152 $\pm$ 0.0166 0.2784 $\pm$ 0.0015
0.6852 $\pm$ 0.0047 5.6055 $\pm$ 0.0153 0.2720 $\pm$ 0.0013
0.7614 $\pm$ 0.0049 5.7146 $\pm$ 0.0162 0.2827 $\pm$ 0.0008
0.7703 $\pm$ 0.0036 5.8232 $\pm$ 0.0197 0.2856 $\pm$ 0.0015
@@ 5 Conclusion
In this paper, we proposed a novel training strategy for diffusion and consistency models, which is based on introducing curriculum learning into Direct Preference Optimization. Furthermore, we integrated DPO into a recent category of efficient diffusion models, known as consistency models. Through comprehensive experiments consisting of three distinct tasks, we demonstrated significant performance gains due to our progressive training based on easy-to-hard image pairs. Aside from presenting numerical results on representative automatic metrics to assess our method and determine its benefits, we also carried out a subjective human evaluation study. The human evaluation study confirmed our observations, namely that curriculum learning brings significant performance gains when integrated into DPO. We thus conclude that our contribution leads to the development of stronger generative models that can accurately capture and synthesize human-preferred images. In future work, we aim to extend and apply Curriculum DPO to LLMs.
Appendix A Supplemental Material
A.1 Detailed Preliminaries
Diffusion models. Diffusion models [10, 19, 25, 27, 49, 59, 63, 62, 69] are a class of generative models trained to reverse a process that progressively inserts Gaussian noise across $T$ steps to the original data samples, transforming them into standard Gaussian noise. Formally, the forward process defined by:
@@ A.1 Detailed Preliminaries
Consistency models. Consistency models [34, 61, 64] are a new class of generative models. These models operate on the idea of training a model to associate each point along a trajectory of the PF-ODE (Eq. (13)) to the trajectory’s initial point, which corresponds to the denoised sample. Such models can either be trained from scratch or through distillation from a pre-trained diffusion model. In our study, we used the distillation method, so we next detail this approach.
$\mathcal{L}_{\scriptsize{\mbox{CD}}}(\phi)=d(f_{\phi}(x_{t_{n+1}},t_{n+1}),f_{\phi^{-}}(\hat{x}_{t_{n}}^{\theta},t_{n})),$ (14)
[…]
$p_{\scriptsize{\mbox{BT}}}(x_{0}^{w}\prec x_{0}^{l}|c)=\sigma(r_{\varphi}(x_{0}^{w},c)-r_{\varphi}(x_{0}^{l},c)),$ (15)
$\mathcal{L}_{\scriptsize{\mbox{BT}}}=-\mathbb{E}_{x_{0}^{w},x_{0}^{l},c}\left(\log{\sigma(r_{\varphi}(x_{0}^{w},c)-r_{\varphi}(x_{0}^{l},c))}\right).$ (16)
After training the reward model $r_{\varphi}(x_{0},c)$, RLHF optimizes a conditional generative model $p_{\theta}(x_{0}|c)$ to maximize the reward $r_{\varphi}(x_{0},c)$ and, at the same time, controls the deviance from a reference model $p_{\scriptsize{\mbox{ref}}}(x_{0},c)$ through a Kullback–Leibler (KL) divergence term:
$\max_{\theta}\mathbb{E}_{c,x_{0}\sim p_{\theta}(x_{0}|c)}r_{\varphi}(x_{0},c)-\beta KL(p_{\theta}(x_{0}|c),p_{\scriptsize{\mbox{ref}}}(x_{0}|c)),$ (17)
where $\beta$ controls the importance of the divergence term.
@@
$p_{\theta}^{*}(x_{0}|c)=\frac{p_{\scriptsize{\mbox{ref}}}(x_{0}|c)\cdot\exp\left(\frac{r(x_{0},c)}{\beta}\right)}{Z(c)},$ (18)
where $Z(c)=\sum_{x_{0}}{p_{\scriptsize{\mbox{ref}}}(x_{0}|c)\cdot\exp\left(\frac{r(x_{0},c)}{\beta}\right)}$ is a normalization constant. Further, from Eq. (18), Rafailov et al. [46] rewrite the reward as:
$r(x_{0},c)=\beta\left(\log\frac{p_{\theta}^{*}(x_{0}|c)}{p_{\scriptsize{\mbox{ref}}}(x_{0}|c)}+\log Z(c)\right).$ (19)
@@
$\mathcal{L}_{\mathrm{\scriptsize{\mbox{DPO}}}}(\theta)=-\mathbb{E}_{x_{0}^{w},x_{0}^{l},c}\left(\log{\sigma\left(\beta\cdot\left(\log{\frac{p_{\theta}(x_{0}^{w}|c)}{p_{\scriptsize{\mbox{ref}}}(x_{0}^{w}|c)}}-\log{\frac{p_{\theta}(x_{0}^{l}|c)}{p_{\scriptsize{\mbox{ref}}}(x_{0}^{l}|c)}}\right)\right)}\right).$ (20)
To grasp the intuition behind $\mathcal{L}_{\mathrm{\scriptsize{\mbox{DPO}}}}$, we can analyze its gradient with respect to $\theta$:
$\frac{\partial\mathcal{L}_{\mathrm{\scriptsize{\mbox{DPO}}}}(\theta)}{\partial\theta}\!=\!-\beta\mathbb{E}_{x_{0}^{w},x_{0}^{l},c}\left[\sigma\left(\hat{r}_{\theta}(x_{0}^{l},c)-\hat{r}_{\theta}(x_{0}^{w},c)\right)\!\cdot\!\left(\frac{\partial\log{p_{\theta}(x_{0}^{w}|c)}}{\partial\theta}-\frac{\partial\log{p_{\theta}(x_{0}^{l}|c)}}{\partial\theta}\right)\!\right],$ (21)
with $\hat{r}_{\theta}(x_{0},c)=\beta\cdot\log{\frac{p_{\theta}(x_{0}|c)}{p_{\scriptsize{\mbox{ref}}}(x_{0}|c)}}$. By analyzing Eq. (21), as discussed in [46], it is evident that the DPO objective enhances the likelihood of favored examples, while diminishing it for the unfavored ones. The magnitude of the update is proportional to the error in $\hat{r}_{\theta}$. Here, the term “error” refers to the degree to which $\hat{r}_{\theta}$ incorrectly prioritizes the sample $x_{0}^{l}$.
A.2 Curriculum DPO for Diffusion Models

### v3

@@ 4 Experiments
Dataset $D_{1}$ [7] Dataset $D_{2}$ [54]
Model Fine-Tuning Strategy Text Aesthetics Human Text Aesthetics Human
Alignment Preference Alignment Preference
LCM - 0.7243±0.0048 6.0490±0.0162 0.2912±0.0021 0.5602±0.0032 5.8038±0.0139 0.2610±0.0016
DDPO [7] 0.7490±0.0036 6.3730±0.0130 0.2952±0.0011 0.5721±0.0043 6.0121±0.0127 0.2803±0,0019
DPO [72] 0.7502±0.0045 6.4741±0.0095 0.2990±0.0010 0.5720±0.0040 6.0430±0.0113 0.2814±0.0023
Curriculum DPO (ours) 0.7548±0.0041 6.6417±0.0083 0.3237±0.0012 0.5812±0.0038 6.1829±0.0128 0.2851±0.0017
SD - 0.6804±0.0052 5.5152±0.0166 0.2784±0.0015 0.5997±0.0067 5.4292±0.0181 0.2646±0.0011
DDPO [7] 0.7629±0.0040 5.8183±0.0129 0.2854±0.0012 0.6024±0.0055 5.6748±0.0162 0.2673±0.0025
DPO [72] 0.7614±0.0049 5.7146±0.0162 0.2827±0.0008 0.6075±0.0057 5.6205±0.0152 0.2672±0.0013
Curriculum DPO (ours) 0.7703±0.0036 5.8232±0.0197 0.2856±0.0015 0.6234±0.0050 5.7060±0.0177 0.2681±0.0019
@@
Table 1: Text alignment, aesthetic and human preference scores on datasets $D_{1}$ and $D_{2}$, obtained by the baseline (pre-trained) LCM and SD models versus the three fine-tuning strategies: DDPO, DPO and Curriculum DPO. The results represent averages over three runs. The best scores are highlighted in bold.
Figure 2: Qualitative results on dataset $D_{1}$, before and after fine-tuning for the text alignment task. The fine-tuning alternatives are: DDPO, DPO and Curriculum DPO. Best viewed in color.
@@
Datasets. We run our experiments on three datasets. The first dataset ($D_{1}$) is created with the procedure of Black et al. [7]. For the text alignment and human preference tasks, the text prompts are based on the subject-verb-object (SVO) pattern, e.g. “a dog riding a bike”. Following Black et al. [7], we use a list of 45 animals and 3 types of activities. For the aesthetics task, the prompts are based on the “a photo of <an object>” template, where the object represents one of the same 45 animal classes. We generate 500 images for each text prompt. For the text alignment and human preference tasks, there are 67,500 (prompt, image) pairs for training, and 6,750 pairs for evaluation. For the aesthetics task, there are 22,500 pairs for training and 2,250 for evaluation.
The second dataset ($D_{2}$) is DrawBench [54], which consists of 200 diverse prompts. As for the first dataset, we generate 500 images per prompt. Therefore, the training set consists of 100,000 (prompt, image) pairs, while the test set comprises 10,000 pairs.
The third dataset ($D_{3}$) is a subset of 150,000 (prompt, image) pairs from Pick-a-Pic [29]. For $D_{3}$, the generated images are already provided. The test set contains 500 prompts.
@@
Generative models. We conduct our experiments by using two pre-trained text-to-image generative models: Stable Diffusion (SD) v1.5 [51] and Latent Consistency Model (LCM) [36]. The LCM checkpoint is trained using consistency distillation from an SD v1.5 checkpoint. To generate images with SD, we use the DDIM sampler with 50 steps. For LCM, we generate data using 8 steps of the Multistep Latent Consistency Sampling [36] procedure. Finally, we consider two different resolutions, $256\times 256$ pixels for SD and $768\times 768$ pixels for LCM, respectively.
@@
Reward models. To assess the alignment between a text prompt and a generated image, we use the cosine similarity to compare the embeddings produced by Sentence-BERT [50] for the original prompt and the caption generated by LLaVA [33]. To evaluate visual appeal (aesthetics), we use the LAION Aesthetics Predictor [57], a linear model applied on CLIP [47] that is trained on human-rated images. Finally, for human preference estimation, we employ HPSv2 [75], a CLIP model fine-tuned on a dataset of image pairs ranked by humans. The reward models are established according to Black et al. [7]. In each scenario, DPO, DDPO and Curriculum DPO use the same reward models, ensuring a fair comparison. We present results with an additional reward model (Phi-3 [1]) in Table 5 from the supplementary.
@@
Training setup. We conduct all experiments on a single A100 GPU, training each model for 10,000 training iterations, with a batch size of 16 and two steps of gradient accumulation. The training time is approximately 2 days for each LCM experiment, while using 64GB of VRAM. The SD experiments take approximately 1 day each, while using 36 GB of VRAM. We employ the AdamW optimizer with a constant learning rate of $3\cdot 10^{-4}$. LoRA is activated in all experiments. For LCM, we utilize trainable low-rank matrices with a dimension of 64 and an equivalent $\alpha$. In the case of SD, we employ a rank of 8 and an $\alpha$ of 32. Following Wallace et al. [72], we set $\beta=5000$ for Diffusion-DPO. For Consistency-DPO, we set $\beta=200$ via grid search over the set $\{50,100,200,300,500\}$. Similarly, we use grid search over the set $\{3,5,7\}$ to determine the number of curriculum batches for Curriculum DPO. The optimal value is $B=5$. The number of training iterations per batch is set to $K=400$ for the first four batches, i.e. $H_{i}=K,\forall i\in\{1,2,3,4\}$. For the last batch, the number of iterations is set to $H_{5}=10000-4\cdot K$. Hence, the total number of training iterations is the same as for the baseline, i.e. $\sum_{i=1}^{B}H_{i}=10000$. Since all experiments involve reward models, the image pairs are sampled based on their preference score difference rather than their ranking difference, thus sidestepping the management of skewed preference score distributions.
@@
Results. For the quantitative assessment, we measure the reward scores after fine-tuning the models on each of the three tasks, corresponding to the three reward models. The results on datasets $D_{1}$ and $D_{2}$ are shown in Table 1 (results on $D_{3}$ are presented in Table 4 from the supplementary). As anticipated, the original models yield the lowest scores, indicating that fine-tuning is generally useful. Both DPO and DDPO exhibit better performance than the baseline, but without a clear winner among the two approaches. In contrast, Curriculum DPO is the most effective method, achieving the best reward score on each task and each dataset. Moreover, the gains of Curriculum DPO are significant in most cases.
@@
We present qualitative results for the text alignment task in Figure 2. Considering the examples from the first column, we observe that Curriculum DPO is the only method which generates the beetle and the bike in a coherent picture. Moreover, in the third column, our method is the only one able to generate the butterfly without additional humans around. Regarding the examples generated with SD, Curriculum DPO exhibits clear improvements over the other methods when generating the goose (last column) and the squirrel riding a bike (sixth column). Qualitative results for the aesthetics and human preference tasks are presented in Appendix 10.
@@
- 2.778 2.276 2.718 3.510
DDPO 2.810 2.983 2.765 3.454
DPO 2.846 2.821 2.782 2.906
Curriculum DPO (ours) 3.440 3.175 3.006 3.664
Table 2: Average ratings of the baseline models and the three fine-tuning strategies based on our human evaluation study. The study was completed by 9 human evaluators. The best average ratings are highlighted in bold. Statistical testing indicates that the voting results are statistically significant, with a p-value below 0.005.
@@
Figure 3: Ablation results obtained by varying the hyperparameter $\beta$ for Consistency-DPO, the number of training iterations per batch $K$ for Curriculum DPO, the number of batches $B$ for Curriculum DPO, and the number of training images per prompt $M$. Fine-tuned LCM models are compared with the pre-trained LCM baseline on the human preference and visual appeal tasks (where scores are given by the HPSv2 reward model and LAION Aesthetics Predictor respectively).
@@
Model LoRA DPO Curriculum DPO Text Alignment Aesthetics Human Preference
LCM 0.7243±0.0048 6.0490±0.0162 0.2912±0.0021
0.7151±0.0032 5.7158±0.0154 0.2881±0.0017
0.7502±0.0045 6.4741±0.0095 0.2990±0.0010
0.7548±0.0041 6.6417±0.0083 0.3237±0.0012
SD 0.6804±0.0052 5.5152±0.0166 0.2784±0.0015
0.6852±0.0047 5.6055±0.0153 0.2720±0.0013
0.7614±0.0049 5.7146±0.0162 0.2827±0.0008
0.7703±0.0036 5.8232±0.0197 0.2856±0.0015
Table 3: Ablation study on the key components of Curriculum DPO. LoRA, DPO and curriculum learning are gradually added to produce Curriculum DPO. The results represent averages over three runs. The best scores are highlighted in bold. Statistical significance testing for DPO versus Curriculum DPO attest that the differences brought by curriculum learning are significant (p-values are below 0.005).
Subjective human evaluation study. We conducted a human evaluation study to evaluate the prompt alignment and aesthetics of the two pre-trained models, SD and LCM, along with their fine-tuned versions based on DDPO, DPO and Curriculum DPO, respectively. We asked the annotators to rate each image with a grade from $1$ to $5$, considering the task-specific evaluation criteria (prompt alignment / aesthetics). For each prompt, the generated images were shuffled and presented in a random order to prevent any form of cheating from the annotators. More details about the annotation process are provided in Appendix 12. We present the average ratings in Table 2. We observe that Curriculum DPO outperforms the other methods by considerable margins, regardless of the task.
Ablation studies. We conduct several ablation studies to explore the impact of various hyperparameters associated with Consistency-DPO and Curriculum DPO. In Figure 3(a), we show various values for $\beta$ and their impact on DPO when applied on consistency models. For values of $\beta$ higher or equal to 200, Consistency-DPO surpasses the pre-trained LCM baseline. In Figure 3(b), we present an ablation on the impact of $K$ (the number of training iterations per curriculum batch) on the human preference scores returned by HPSv2. We observe better results for $K=300$ and $K=400$, although Curriculum DPO surpasses the baseline LCM for all values of $K$. In Figure 3(c), we present the HPSv2 scores for various choices of the number of batches, $B$. Once again, Curriculum DPO consistently surpasses the pre-trained LCM, regardless of the number of training batches. In Figure 3(d), we vary the number of generated images per prompt from $M=5$ to $M=500$. For reference, we include the results of DPO and DDPO based on $M=500$ images per prompt. Remarkably, Curriculum DPO is able to obtain similar scores to DPO and DDPO, while using 10$\times$ less training images. Moreover, Curriculum DPO outperforms the baseline LCM regardless of the number of training images.
Next, we present a detailed ablation study of the key components of Curriculum DPO in Table 3. This analysis showcases the impact of the LoRA fine-tuning technique, the DPO objective, and the curriculum strategy. Interestingly, applying LoRA to LCM degrades performance on all three tasks. However, combining LoRA with DPO or Curriculum DPO leads to significantly better results. Curriculum DPO consistently outperforms its ablated version, justifying the proposed design.
@@ 5 Conclusion
In this paper, we proposed a novel training strategy for diffusion and consistency models, which is based on introducing curriculum learning into Direct Preference Optimization. Furthermore, we integrated DPO into a recent category of efficient diffusion models, known as consistency models. Through comprehensive experiments consisting of three distinct tasks, we demonstrated significant performance gains due to our progressive training based on easy-to-hard image pairs. Aside from presenting numerical results on representative automatic metrics to assess our method and determine its benefits, we also carried out a subjective human evaluation study. The human evaluation study confirmed our observations, namely that curriculum learning brings significant performance gains when integrated into DPO. We thus conclude that our contribution leads to the development of stronger generative models that can accurately capture and synthesize human-preferred images.
Acknowledgments. This work was supported by a grant of the Ministry of Research, Innovation and Digitization, CCCDI - UEFISCDI, project number PN-IV-P6-6.3-SOL-2024-2-0227, within PNCDI IV.
Supplementary Material
6 Detailed Preliminaries
Diffusion models. Diffusion models [11, 20, 26, 28, 51, 61, 65, 64, 71] are a class of generative models trained to reverse a process that progressively inserts Gaussian noise across $T$ steps into the original data samples, transforming them into standard Gaussian noise. Formally, the forward process defined by:
@@ 6 Detailed Preliminaries
Consistency models. Consistency models [36, 63, 66] are a new class of generative models. These models operate on the idea of training a model to associate each point along a trajectory of the PF-ODE (Eq. (13)) to the trajectory’s initial point, which corresponds to the denoised sample. Such models can either be trained from scratch or through distillation from a pre-trained diffusion model. In our study, we employ the distillation method, so we next detail this approach.
$\mathcal{L}_{\mbox{\scriptsize{CD}}}(\phi)=d(f_{\phi}(x_{t_{n+1}},t_{n+1}),f_{\phi^{-}}(\hat{x}_{t_{n}}^{\theta},t_{n})),$ (14)
[…]
$p_{\mbox{\scriptsize{BT}}}(x_{0}^{w}\prec x_{0}^{l}|c)=\sigma(r_{\varphi}(x_{0}^{w},c)-r_{\varphi}(x_{0}^{l},c)),$ (15)
$\mathcal{L}_{\mbox{\scriptsize{BT}}}=-\mathbb{E}_{x_{0}^{w},x_{0}^{l},c}\left[\log{\sigma(r_{\varphi}(x_{0}^{w},c)-r_{\varphi}(x_{0}^{l},c))}\right].$ (16)
After training the reward model $r_{\varphi}(x_{0},c)$, RLHF optimizes a conditional generative model $p_{\theta}(x_{0}|c)$ to maximize the reward $r_{\varphi}(x_{0},c)$ and, at the same time, controls the deviance from a reference model $p_{\mbox{\scriptsize{ref}}}(x_{0},c)$ through a Kullback–Leibler (KL) divergence term:
$\max_{\theta}\mathbb{E}_{c,x_{0}\sim p_{\theta}(x_{0}|c)}\left[r_{\varphi}(x_{0},c)\!-\!\beta\mbox{KL}(p_{\theta}(x_{0}|c),p_{\mbox{\scriptsize{ref}}}(x_{0}|c))\right]\!,$ (17)
where $\beta$ controls the importance of the divergence term.
@@
$p_{\theta}^{*}(x_{0}|c)=\frac{p_{\mbox{\scriptsize{ref}}}(x_{0}|c)\cdot\exp\left(\frac{r(x_{0},c)}{\beta}\right)}{Z(c)},$ (18)
where $Z(c)=\sum_{x_{0}}{p_{\mbox{\scriptsize{ref}}}(x_{0}|c)\cdot\exp\left(\frac{r(x_{0},c)}{\beta}\right)}$ is a normalization constant. Further, from Eq. (18), Rafailov et al. [48] rewrite the reward as:
$r(x_{0},c)=\beta\left(\log\frac{p_{\theta}^{*}(x_{0}|c)}{p_{\mbox{\scriptsize{ref}}}(x_{0}|c)}+\log Z(c)\right).$ (19)
@@
$\begin{split}\mathcal{L}_{\mathrm{\mbox{\scriptsize{DPO}}}}(\theta)\!\!=\!\!-\!\mathbb{E}_{x_{0}^{w}\!,x_{0}^{l},c}\!\left[\!\log\!\sigma\!\bigg{(}\!\!\beta\!\!\left(\!\!\log\!{\frac{p_{\theta}(x_{0}^{w}\!|c)}{p_{\mbox{\scriptsize{ref}}}(x_{0}^{w}\!|c)}}\!-\!\log\!{\frac{p_{\theta}(x_{0}^{l}\!|c)}{p_{\mbox{\scriptsize{ref}}}(x_{0}^{l}\!|c)}}\!\!\right)\!\!\!\!\bigg{)}\!\!\right]\!\!,\end{split}$ (20)
To grasp the intuition behind $\mathcal{L}_{\mathrm{\mbox{\scriptsize{DPO}}}}$, we can analyze its gradient with respect to $\theta$:
$\begin{split}\frac{\partial\mathcal{L}_{\mathrm{\mbox{\scriptsize{DPO}}}}(\theta)}{\partial\theta}\!=\!&-\beta\mathbb{E}_{x_{0}^{w},x_{0}^{l},c}\Bigg{[}\sigma\left(\hat{r}_{\theta}(x_{0}^{l},c)-\hat{r}_{\theta}(x_{0}^{w},c)\right)\!\cdot\!\\ &\left(\frac{\partial\log{p_{\theta}(x_{0}^{w}|c)}}{\partial\theta}-\frac{\partial\log{p_{\theta}(x_{0}^{l}|c)}}{\partial\theta}\right)\!\Bigg{]},\end{split}$ (21)
with $\hat{r}_{\theta}(x_{0},c)=\beta\cdot\log{\frac{p_{\theta}(x_{0}|c)}{p_{\mbox{\scriptsize{ref}}}(x_{0}|c)}}$. By analyzing Eq. (21), as discussed in [48], it is evident that the DPO objective enhances the likelihood of favored examples, while diminishing it for the unfavored ones. The magnitude of the update is proportional to the error in $\hat{r}_{\theta}$. Here, the term “error” refers to the degree to which $\hat{r}_{\theta}$ incorrectly prioritizes the sample $x_{0}^{l}$.
Figure 4: Qualitative results before and after fine-tuning for the text alignment task on DrawBench. The fine-tuning methods are: DDPO, DPO, Naive DPO and Curriculum DPO. Best viewed in color.
7 Curriculum DPO for Diffusion Models